OpenStudy (anonymous):

Which mixed number falls between 4 3/8 and 5 5/12 ? A) 4 1/4 B) 5 2/3 C) 4 1/2 D) 5 5/6

OpenStudy (anonymous):

The easiest way to do this is to turn all of the fractions into decimals and just put them on a number line

OpenStudy (anonymous):

Ok thanks im gonna do that now

OpenStudy (anonymous):

i got D is that right?

hero (hero):

It's not D

hero (hero):

There's a much easier way to do these without a number line

OpenStudy (anonymous):

im super confused...

hero (hero):

I understand. You need someone to explain it you.

hero (hero):

Basically, there's a simple method for this, which only necessitates paying attention to the fractions.

hero (hero):

There are basically three types of fractions: Fractions that are either zero or close to zero Fractions that are either half or close to half Fractions that are either one or close to one Obviously 0 < 1/2 < 1

hero (hero):

If you have something like 1/3 , 1/4 , 1/5 those are close to zero

hero (hero):

If you have something like 4/5, 5/6, 7/8, those are close to one

hero (hero):

If you have something like 3/8, 4/10, 5/12, those are close to half

hero (hero):

obviously, the fractions that are close to zero will be less than those that are close to half, which in turn, are less than those close to one.

hero (hero):

In the case of your problem above we have 4 3/8 and 5 5/12

hero (hero):

Therefore, we are only interested in numbers that are greater than 4 3/8 but less than 5 5/12

hero (hero):

For the first possibility, 4 1/4... 1/4 is close to zero because if you wrote out the fractions in order: 0/4, 1/4, 2/4, 3/4.... 0/4 = 0 and 1/4 is obviously close to that, but 3/8 is close to half since 4/8 = 1/2

hero (hero):

If you wrote out the eighths in order you would have 0/8, 1/8, 2/8, `3/8, 4/8`, 5/8, 6/8, 7/8, 8/8

hero (hero):

If 1/4 is close to zero and 3/8 is close to 1/2, then 1/4 must be less than 3/8

hero (hero):

Therefore, we can conclude that 4 1/4 is not greater than 4 3/8...so that is not a possible choice.

OpenStudy (anonymous):

im still here lol just busy taking a test